# A classification of Fourier summation formulas and crystalline measures

### v1

@@ 3. Applications
where $E^{*}(z)=\overline{E(\overline{z})}$. In this case we always write $E=A-iB$, where $A$ and $B$ are real entire functions [^9] Entire functions which attain only real values on the real line. defined by $A=(E^{*}+E)/2$ and $B=(E^{*}-E)/(2i)$, and we always write $\Theta=E^{*}/E$. In particular we have the following identity
@@
Let $E=A-iB$ be of Hermite-Biehler class such that $A/B\in\mathrm{AP}(\mathbb{C}^{+})$ and $\lambda\in\mathbb{R}\mapsto\mathbb{E}(A/B)(\lambda)$ is locally summable. Let
@@
Then $(\mu,a)$ is a real-antipodal $\mathrm{FS}$-pair such that $\mu\geq 0$ and $\mu$ has locally finite support. Conversely, if $(\mu,a)$ is a real-antipodal $\mathrm{FS}$-pair such that $\mu\geq 0$, $\operatorname{\mathrm{supp}}(\mu)$ is locally finite and $a(\cdot)$ has exponential growth, then $(\mu,a)$ has to be built from the construction above.
@@
Observe that by Lemma 11(iii),(iv),(v) and identity (7), $A/B\in\mathrm{AP}(\mathbb{C}^{+})$ if and only if $\Theta=E^{*}/E\in\mathrm{AP}(\mathbb{C}^{+})$ and $A/B$ has locally finite spectrum if and only if $\Theta$ has. Note also that if $E,E^{*}\in\mathrm{AP}(\mathbb{C}^{+})$ then $\Theta\in\mathrm{AP}(\mathbb{C}^{+})$. In particular, if $E$ is a trigonometric polynomial of Hermite-Biehler class, then $iA/B$ belongs to AP(+)normal-$\mathrm{AP}(\mathbb{C}^{+})$ and has locally finite spectrum, hence Theorem 3 applies and the pair $(\mu,a)$ is crystalline ($(\mu,a)\in\rm CM$, see Appendix).
@@
Theorem 3 generalizes a construction of Kurasov & Sarnak [21]. They construct a crystalline $\mathrm{FS}$-pair $(\mu,a)\in\rm CM$ by letting
@@
We obtain that $E=Q^{\prime}-iQ$ is a trigonometric polynomial of Hermite-Biehler class by a routine calculation. By the previous remark, Theorem 3 applies. It is also easy to see that the zeros of $Q$ coincide with the points $\gamma\in\mathbb{R}$ such that $\varphi_{E}(\gamma)\equiv 0\,({\rm mod}\,\pi)$ and $\varphi^{\prime}_{E}(\gamma)^{-1}=\mathop{\mathrm{Res}}_{z=\gamma}\,(A/B)=m(\gamma)$. Thus, by Theorem 3 (with $a(\cdot)$ as in (8)), both $\mu$ and $a(\cdot)$ have locally finite support, and so $(\mu,a)$ is crystalline ($(\mu,a)\in\mathrm{RRTP}$, see Appendix).
@@
[…]
@@
Let $f=iA/B$. Since $E=A-iB$ is Hermite-Biehler, we obtain that ${\rm Re}\,f>0$ in $\mathbb{C}^{+}$, and so $|e^{-f}|<1$ in $\mathbb{C}^{+}$, hence $e^{f}$ is of bounded type. Lemma 11(iv) shows that $\Theta=\frac{1-f}{1+f}\in\mathrm{AP}(\mathbb{C}^{+})$ and that $|\Theta|<1-c_{\varepsilon}$ for ${\rm Im}\,z>\varepsilon$, so $f=\frac{1+\Theta}{1-\Theta}$ is bounded for ${\rm Im}\,z>\varepsilon$, for any $\varepsilon>0$. In particular $\lim_{y\to\infty}{\rm Re}\,f(iy)/y=0$. Moreover, $\mathbb{E}f(\lambda)$ is locally summable by hypothesis. We conclude that Theorem 2 applies. Poisson representation [9, Thm. 4 & Prob. 89] for $f$ implies that $\mu$ has the form given by (8). Conversely, if $(\mu,a)$ is a real-antipodal $\mathrm{FS}$-pair such that $\mu\geq 0$, $\operatorname{\mathrm{supp}}(\mu)$ is locally finite and $a(\cdot)$ has exponential growth, then we can apply Theorem 4 to obtain that (8) holds for some Hermite-Biehler function $E=A-iB$ such that $A(z+ic)/B(z+ic)$ belongs to $\mathrm{AP}(\mathbb{C}^{+})$ for some $c>0$. However, Lemma 11 also shows that $\Theta(z+ic)=E^{*}(z+ic)/E(z+ic)$ belongs to $\mathrm{AP}(\mathbb{C}^{+})$, and since $|\Theta|<1$ in $\mathbb{C}^{+}$, Lemma 10 shows that $\Theta\in\mathrm{AP}(\mathbb{C}^{+})$ and, by Lemma 11 again we obtain that $A/B\in\mathrm{AP}(\mathbb{C}^{+})$. Since $a(\lambda)$ is locally summable and $a(\lambda)=\mathbb{E}f(\lambda)$ for $\lambda>0$, then $\mathbb{E}f(\lambda)$ is locally summable. ∎
@@
Let $(\mu,a)$ be an $\mathrm{FS}$-pair such that $\mu$ has locally finite support, $\deg(\mu)\leq 2$ and $a(\cdot)$ has exponential growth. Then there exists four Hermite-Biehler functions $E_{j}=A_{j}-iB_{j}$ and numbers $p_{j}\in\{0,1\}$, for $j=1,2,3,4$, such that if we let $f=ip_{1}A_{1}/B_{1}-ip_{2}A_{2}/B_{2}$ and $g=ip_{3}A_{3}/B_{3}-ip_{4}A_{4}/B_{4}$ then $f(\cdot+ic)$ and $g(\cdot+ic)$ belong to AP(+)normal-$\mathrm{AP}(\mathbb{C}^{+})$ for some $c>0$, $\mathbb{E}f(\lambda)$ and $\mathbb{E}g(\lambda)$ are locally summable, and:
$\frac{1}{2\pi}{\mu}$ $=p_{1}\sum_{\,\varphi_{1}(\gamma)\equiv 0\,({\rm mod}\,\pi)}\frac{1}{\varphi_{1}^{\prime}(\gamma)}{{\boldsymbol{\delta}}}_{\gamma}-p_{2}\sum_{\,\varphi_{2}(\gamma)\equiv 0\,({\rm mod}\,\pi)}\frac{1}{\varphi_{2}^{\prime}(\gamma)}{{\boldsymbol{\delta}}}_{\gamma}$
$\ \ -ip_{3}\sum_{\,\varphi_{3}(\gamma)\equiv 0\,({\rm mod}\,\pi)}\frac{1}{\varphi_{3}^{\prime}(\gamma)}{{\boldsymbol{\delta}}}_{\gamma}+ip_{4}\sum_{\,\varphi_{4}(\gamma)\equiv 0\,({\rm mod}\,\pi)}\frac{1}{\varphi_{4}^{\prime}(\gamma)}{{\boldsymbol{\delta}}}_{\gamma};$
[…]
$\tfrac{1}{2}a(0)$ $={\rm Re}\,\mathbb{E}f(0)-i{\rm Re}\,\mathbb{E}g(0).$
@@
for $j=1,2$, where $\mu_{j}=\sum_{\gamma\in\Lambda}r_{j}(\gamma){\boldsymbol{\delta}}_{\gamma}$. Observe that each $f_{j}$ is meromorphic in $\mathbb{C}$ with only simple poles (possibly) at $\Lambda$. Since $\mu_{j}\geq 0$ we have $if_{j}(\mathbb{C}^{+})\subset\mathbb{C}^{+}\cup\{0\}$, and we can then apply a classical result [26, p. 308, Thm. 1] to deduce this happens exactly when there are two real entire functions $A_{j}$ and $B_{j}$ with only real zeros that interlace and such that $if_{j}=-p_{j}A_{j}/B_{j}$, with $p_{j}\in\{0,1\}$ to account for the case when $\mu_{j}=0$ for some $j=0,1$. Since ${\rm Re}\,iA_{j}/B_{j}>0$, we deduce that $E_{j}=A_{j}-iB_{j}$ is of Hermite-Biehler class, and so, Poisson representation shows that
@@
In what follows we say that an entire function $E$ is of Pólya class if $E$ has no zeros in $\mathbb{C}^{+}$, $|E^{*}/E|\leq 1$ in $\mathbb{C}^{+}$ and $y\mapsto|E(x+iy)|$ is nondecreasing for $y>0$. A classical theorem [9, Thm. 7] shows that a function $E$ is of Pólya class if and only if
[…]
@@
The Pólya class can be also defined as functions that are uniform limits in compact sets of polynomials with no zeros in $\mathbb{C}^{+}$ (see [9, Problem 12]). Moreover, any function of Pólya class has order at most $2$ ([9, Problem 10]).
@@
Let $(\mu,a)$ be an $\mathrm{FS}$-pair. Assume that
@@
    There is $b>0$ such that $\operatorname{\mathrm{supp}}(a)\cap(0,b)=\emptyset$ and $a(\cdot)$ has the decay (5).
Then there exists an entire function $E=A-iB$ of Hermite-Biehler class and of finite exponential type such that $E,E^{*}\in\mathrm{AP}(\mathbb{C}^{+})$ (and so $A/B\in\mathrm{AP}(\mathbb{C}^{+})$), ${\rm spec}(E)$ is bounded and (8) holds. If in addition $\operatorname{\mathrm{supp}}(a)$ is locally finite, then $E$ is a trigonometric polynomial and Remark 5 applies, and so $(\mu,a)\in\mathrm{RRTP}$ (see Appendix).
@@
Step 1. We first apply Theorem 3 to conclude there is an Hermite-Biehler function $E=A-iB$ such that $A/B\in\mathrm{AP}(\mathbb{C}^{+})$ and (8) holds. We can assume that $E$ has no real zeros as these can be removed (with a Weierstrass product) without altering $A/B$ or the fact that $E$ is Hermite-Biehler. In particular, $\operatorname{\mathrm{supp}}(\mu)=\{\gamma\in\mathbb{R}:\varphi(\gamma)\equiv 0\,({\rm mod}\,\pi)\}=\text{Zeros}(B)$, where $\varphi$ is the phase function associated with $E$, and these zeros are all simple. The zeros of $A$ are also simple and interlace those of $B$. Since $\mu$ is bounded from below on its support and that $a(\cdot)$ has the decay (5), we can apply Remark 3 to obtain
@@
for some $10<c<$ 1, whenever $B(\gamma)=0$. In particular, we have $\sum_{B(\gamma)=0}\frac{1}{1+\gamma^{2}}<\infty$ and, since the zeros of $A$ interlace those of $B$, we also have $\sum_{A(s)=0}\frac{1}{1+s^{2}}<\infty$. We claim that we can assume $E$ is has order at most $1$. Indeed, define the canonical products
@@
for some $p,q\in\{0,1\}$, where $b\in\mathbb{R}$ is chosen so that $(A(z)/B(z))/(\mathcal{A}(z)/\mathcal{B}(z))=1+O(z)$ for $z\to 0$. Hence both $\mathcal{A}$ and $\mathcal{B}$ have order at most $1$ by a classical result of Borel [26, Thm. 6, p. 16]. Since ${\rm Re}\,iA/B\geq 0$ in $\mathbb{C}^{+}$ and, a straitfoward computation, shows that ${\rm Re}\,i\mathcal{A}/\mathcal{B}\geq 0$ in $\mathbb{C}^{+}$, we conclude that both $A/B$ and $\mathcal{A}/\mathcal{B}$ are of bounded type and so $F=(A/B)/(\mathcal{A}/\mathcal{B})$ also is and has no zeros. Since $F=F^{*}$, a classical result of Krein [20] (see also [27, Thm. 1, p. 115]) shows that $F$ is of exponential type, and so $F(z)=e^{hz}$, for some $h\in\mathbb{R}$. Since $F$ is of bounded type we have $h=0$. We can then define $\widetilde{E}=\mathcal{A}-i\mathcal{B}$ to finish the claim.
Step 2. Let $\{\overline{z}_{n}=x_{n}-iy_{n}\}$ be the zeros of $E$ with $y_{n}>0$. We claim $\sup_{n}y_{n}<\infty$ and that $E$ is of Póyla class. Indeed, since $f=iA/B\in\mathrm{AP}(\mathbb{C}^{+})$, by Remark 4 we have $\Theta=E^{*}/E=(1-f)/(1+f)\in\mathrm{AP}(\mathbb{C}^{+})$. Since ${\rm spec}(f)\subset[0,\infty)$, a simple computation using Lemma 11(i) shows $\mathbb{E}\Theta(0)=\frac{1-\mathbb{E}f(0)}{1+\mathbb{E}f(0)}$. However, $02{\rm Re}\,\mathbb{E}f(0)=a(0)>$ 0 (because $\mu\geq 0$), thus $|\mathbb{E}\Theta(0)|<1$. Assume now $\mathbb{E}\Theta(0)\neq 0$. Then Lemma 11(i) shows straightforwardly that $\Theta(x+iy)$ has no zeros for large $y>0$. Assume otherwise that $\mathbb{E}\Theta(0)=0$, that is, $\mathbb{E}f(0)=1$. Since $a(\lambda)=\mathbb{E}f(\lambda)$ for $\lambda>0$ we deduce that ${\rm spec}(f)\subset\{0\}\cup[b,\infty)$ where $b=\inf\{\lambda>0:a(\lambda)\neq 0\}>0$ by assumption. Since $f$ has no zeros in $\mathbb{C}^{+}$, we can then apply Lemma 11(ii) to conclude that $b\in{\rm spec}(f)$. Writing $f=1-2g$, with ${\rm spec}(g)\subset[b,\infty)$ and $\mathbb{E}g(b)\neq 0$, we can apply Lemma 11(i) to conclude that $|g(x+iy)|<1$ for large $y>0$, and so
$\Theta(z)=\frac{g(z)}{1-g(z)}=\sum_{n\geq 1}2^{-n}g(z)^{n},$
@@
Enumerate the $\text{Zeros}(B)=\{\gamma_{n}\}$ and let $c>0$ be so small that $\gamma_{n+1}-\gamma_{n}\geq c$ for all $n$. Let $100<\varepsilon<c/$ 10 to be chosen later. Observe first that if ${\rm dist}(x,\{\gamma_{n}\})=|x-\gamma|\geq\varepsilon$ (with $\gamma=\gamma_{n_{0}}$) then $|x-\gamma_{n_{0}+k}|\geq\varepsilon+|k|c/2$, and so

### v3

@@ 3. Applications
where $E^{*}(z)=\overline{E(\overline{z})}$. In this case we always write E=AiBE=A-iB, where $A$ and $B$ are real entire functions [^10] Entire functions which attain only real values on the real line. defined by $A=(E^{*}+E)/2$ and $B=(E^{*}-E)/(2i)$, and we always write $\Theta=E^{*}/E$. In particular we have the following identity
@@
Let E=AiBE=A-iB be of Hermite-Biehler class such that $A/B\in\mathrm{AP}(\mathbb{C}^{+})$ and $\lambda\in\mathbb{R}\mapsto\mathbb{E}(A/B)(\lambda)$ is locally summable. Let
@@
Then $(\mu,a)$ is a real-antipodal $\mathrm{FS}$-pair such that $\mu\geq 0$ and $\mu$ has locally finite support [^12] A set $S\subset\mathbb{R}$ is locally finite if $(S\cap(a,b))<\infty$ for any $a<b$.. Conversely, if $(\mu,a)$ is a real-antipodal $\mathrm{FS}$-pair such that $\mu\geq 0$, $\operatorname{\mathrm{supp}}(\mu)$ is locally finite and $a(\cdot)$ has exponential growth, then $(\mu,a)$ has to be built from the construction above.
@@
Observe that by Lemma 12(iii),(iv),(v) and identity (6), $A/B\in\mathrm{AP}(\mathbb{C}^{+})$ if and only if $\Theta=E^{*}/E\in\mathrm{AP}(\mathbb{C}^{+})$ and $A/B$ has locally finite spectrum if and only if $\Theta$ has. Note also that if $E,E^{*}\in\mathrm{AP}(\mathbb{C}^{+})$ then $\Theta\in\mathrm{AP}(\mathbb{C}^{+})$. In particular, if $E$ is a trigonometric polynomial of Hermite-Biehler class, then $iA/B$ belongs to $\mathrm{AP}(\mathbb{C}^{+})$ and has locally finite spectrum, hence Theorem 3 applies and the pair $(\mu,a)$ is crystalline.
@@
Theorem 3 generalizes a construction of Kurasov & Sarnak [26]. They construct a crystalline $\mathrm{FS}$-pair $(\mu,a)$ by letting
@@
We obtain that $E=Q^{\prime}-iQ$ is a trigonometric polynomial of Hermite-Biehler class by a routine calculation. By the previous remark, Theorem 3 applies. It is also easy to see that the zeros of $Q$ coincide with the points $\gamma\in\mathbb{R}$ such that $\varphi_{E}(\gamma)\equiv 0\,({\rm mod}\,\pi)$ and $\varphi^{\prime}_{E}(\gamma)^{-1}=\mathop{\mathrm{Res}}_{z=\gamma}\,(A/B)=m(\gamma)$. Thus, by Theorem 3 (with $a(\cdot)$ as in (7)), both $\mu$ and $a(\cdot)$ have locally finite support, and so $(\mu,a)$ is crystalline. A simple example is given by the Hermite-Biehler trigonometric polynomial
$E(z)=\tfrac{1}{2}({(\pi+1)}e^{-\pi iz}+e^{iz}+{(\pi-1)}e^{\pi iz})=\pi\cos(\pi z)+\tfrac{1}{2}\cos(z)-i(\sin(\pi z)+\tfrac{1}{2}\sin(z)),$
which produces the crystalline measure $\mu=\sum_{\sin(\pi\gamma)+\tfrac{1}{2}\sin(\gamma)=0}{\boldsymbol{\delta}}_{\gamma}$.
@@
We now explain how one can prove the main result of Ulanovskii & Oleveskii [40] using Theorem 3, but under milder decay conditions. We claim that if $(\mu,a)$ is an $\mathrm{FS}$-pair such $\mu\geq 0$ is $\mathbb{N}$-valued, $a(\cdot)$ has locally finite support and exponential growth then the Hermite-Biehler function E=AiBE=A-iB given by Theorem 3 is a trigonometric polynomial, and so the construction from Remark 4 applies. Indeed, since $\deg(\mu)\leq 2$ one can take $\mathcal{B}$ to have the same zeros of $B$, but with multiplicities given by the weights of $\mu$. We can take such $\mathcal{B}$ with $\text{order}(\mathcal{B})\leq 1$ by [32, Thm. 6, p. 16]. Using (3) we conclude that ${\rm Re}\,i\mathcal{B}^{\prime}/\mathcal{B}={\rm Re}\,A/B$ in $\mathbb{C}^{+}$, and so Poisson representation shows that $A/B=\mathcal{B}^{\prime}/\mathcal{B}+h$ for some $h\in\mathbb{R}$, hence we can assume that $A=B^{\prime}$. Since $a(\cdot)$ has locally finite support, then $B^{\prime}/B$ has locally finite spectrum contained in $0,\infty)$. Lemma 13 shows that $B\in\mathrm{AP}(\mathbb{C}^{+})$ and it has locally finite spectrum bounded from below. The class $\mathrm{AP}(\mathbb{C}^{+})$ is closed under differentiation, and so $B^{\prime}\in\mathrm{AP}(\mathbb{C}^{+})$, and therefore, $E\in\mathrm{AP}(\mathbb{C}^{+})$, both with locally finite spectrum bounded from below. Note that $B^{\prime}$ also has order $\leq 1$, and so $E$ has order $\leq 1$. We can then apply Lemma 17 to conclude that $E$ is a trigonometric polynomial.
@@
Let $f=iA/B$. Since E=AiBE=A-iB is Hermite-Biehler, we obtain that ${\rm Re}\,f>0$ in $\mathbb{C}^{+}$, and so $|e^{-f}|<1$ in $\mathbb{C}^{+}$, hence $e^{f}$ is of bounded type. Moreover, $\mathbb{E}f(\lambda)$ is locally summable by hypothesis. We conclude that Theorem 2 applies. Poisson representation [13, Thm. 4 & Prob. 89] for $f$ implies that $\mu$ has the form given by (7). Conversely, if $(\mu,a)$ is a real-antipodal $\mathrm{FS}$-pair such that $\mu\geq 0$, $\operatorname{\mathrm{supp}}(\mu)$ is locally finite and $a(\cdot)$ has exponential growth, then we can apply Theorem 4 to obtain that (7) holds for some Hermite-Biehler function E=AiBE=A-iB such that $A(z+ic)/B(z+ic)$ belongs to $\mathrm{AP}(\mathbb{C}^{+})$ for some $c>0$. However, Lemma 12 also shows that $\Theta(z+ic)=E^{*}(z+ic)/E(z+ic)$ belongs to $\mathrm{AP}(\mathbb{C}^{+})$, and since $|\Theta|<1$ in $\mathbb{C}^{+}$, Lemma 11 shows that $\Theta\in\mathrm{AP}(\mathbb{C}^{+})$ and, by Lemma 12 again we obtain that $A/B\in\mathrm{AP}(\mathbb{C}^{+})$. Since $a(\lambda)$ is locally summable and $a(\lambda)=\mathbb{E}f(\lambda)$ for $\lambda>0$, then $\mathbb{E}f(\lambda)$ is locally summable. ∎
@@
Let $(\mu,a)$ be an $\mathrm{FS}$-pair such that $\mu$ has locally finite support, $\deg(\mu)\leq 2$ and $a(\cdot)$ has exponential growth. Then there exists four Hermite-Biehler functions $E_{j}=A_{j}-iB_{j}$ and numbers $p_{j}\in\{0,1\}$, for $j=1,2,3,4$, such that if we let $f=ip_{1}A_{1}/B_{1}-ip_{2}A_{2}/B_{2}$ and $g=ip_{3}A_{3}/B_{3}-ip_{4}A_{4}/B_{4}$ then $f(\cdot+ic)$ and $g(\cdot+ic)$ belong to $\mathrm{AP}(\mathbb{C}^{+})$ for some $c>0$, $\mathbb{E}f(\lambda)$ and $\mathbb{E}g(\lambda)$ are locally summable, and:
$\begin{split}\frac{1}{2\pi}{\mu}&=p_{1}\sum_{\,\varphi_{1}(\gamma)\equiv 0\,({\rm mod}\,\pi)}\frac{1}{\varphi_{1}^{\prime}(\gamma)}{{\boldsymbol{\delta}}}_{\gamma}-p_{2}\sum_{\,\varphi_{2}(\gamma)\equiv 0\,({\rm mod}\,\pi)}\frac{1}{\varphi_{2}^{\prime}(\gamma)}{{\boldsymbol{\delta}}}_{\gamma}\\ &\ \ -ip_{3}\sum_{\,\varphi_{3}(\gamma)\equiv 0\,({\rm mod}\,\pi)}\frac{1}{\varphi_{3}^{\prime}(\gamma)}{{\boldsymbol{\delta}}}_{\gamma}+ip_{4}\sum_{\,\varphi_{4}(\gamma)\equiv 0\,({\rm mod}\,\pi)}\frac{1}{\varphi_{4}^{\prime}(\gamma)}{{\boldsymbol{\delta}}}_{\gamma};\\ a(\lambda)&=\mathbb{E}f(\lambda)-i\mathbb{E}g(\lambda)\quad\text{and}\quad a(-\lambda)=\overline{\mathbb{E}f(\lambda)+i\mathbb{E}g(\lambda)}\quad(\text{for }\lambda>0);\\ \tfrac{1}{2}a(0)&={\rm Re}\,\mathbb{E}f(0)-i{\rm Re}\,\mathbb{E}g(0),\end{split}$ (8)
@@
for $j=1,2$, where $\mu_{j}=\sum_{\gamma\in\Lambda}r_{j}(\gamma){\boldsymbol{\delta}}_{\gamma}$. Observe that each $f_{j}$ is meromorphic in $\mathbb{C}$ with only simple poles (possibly) at $\Lambda$. Since $\mu_{j}\geq 0$ we have $if_{j}(\mathbb{C}^{+})\subset\mathbb{C}^{+}\cup\{0\}$, and we can then apply a classical result [32, p. 308, Thm. 1] to deduce this happens exactly when there are two real entire functions $A_{j}$ and $B_{j}$ with only real zeros that interlace and such that $if_{j}=-p_{j}A_{j}/B_{j}$, with $p_{j}\in\{0,1\}$ to account for the case when $\mu_{j}=0$ for some $j=0,1$. Since ${\rm Re}\,iA_{j}/B_{j}>0$, we deduce that $E_{j}=A_{j}-iB_{j}$ is of Hermite-Biehler class, and so, Poisson representation shows that
@@
Remark 6.
We note that it follows from Theorems 1 and 2 that Theorem 4 is sharp. That is, if we are given four Hermite-Biehler functions $E_{j}=A_{j}-iB_{j}$ and numbers $p_{j}\in\{0,1\}$, for $j=1,2,3,4$, such that $f=ip_{1}A_{1}/B_{1}-ip_{2}A_{2}/B_{2}$ and $g=ip_{3}A_{3}/B_{3}-ip_{4}A_{4}/B_{4}$ belong to $\text{AP}(\mathbb{C}+ic)$ for some $c>0$, then the pair $(\mu,a)$ defined by (8) is an $\mathrm{FS}$-pair. A crucial instantiation of Theorem 4 is the difference of two prime summation formula of Guinand, explained of Section 8.5. It can be obtained from our results in the following way. Let $\chi_{1}$ and $\chi_{2}$ be two primitive and even Dirichlet characters of distinct modulus $N_{1}$ and $N_{2}$, and let
$L(\chi_{j},s)=\sum_{n\geq 1}\chi_{j}(n)n^{-s}$
be their associated L-functions for $j=1,2$ and ${\rm Re}\,s>1$. One can attach to each of them the function
$B_{j}(z)=\widehat{\chi}_{j}(1)^{-1/2}(N_{j}/\pi)^{-iz/2}\Gamma(1/4-iz/2)L(\chi_{j},1/2-iz)$
for complex $z$, where $\widehat{\chi}_{j}(1)=N_{j}^{-1/2}\sum_{n=1}^{N}\chi(j)e^{2\pi in/N}\in\{\pm 1,\pm 1\}$. Thanks to their functional equations, it can be shown that $B_{j}$ is real on the real line, this is, $\overline{B_{j}(\overline{z})}=B_{j}(z)$, and that each $B_{j}$ is an entire function of order $\leq 1$. Assuming the Riemann Hypothesis for both $L$-functions, each $B_{j}$ have only real zeros and those coincide (with multiplicity) with the ordinates of the nontrivial zeros of each $L$-function. We now let $E_{j}=A_{j}-iB_{j}$, with $A_{j}=B^{\prime}_{j}$. Hadamard’s factorization shows that ${\rm Re}\,iA_{j}/B_{j}\geq 0$ in $\mathbb{C}^{+}$ and so each $E_{j}$ is a Hermite-Biehler function. Now observe that
$i\frac{A_{1}}{B_{1}}-i\frac{A_{2}}{B_{2}}$ $=\frac{1}{2}\log(N_{1}/N_{2})+\frac{L^{\prime}(\chi_{1},1/2-iz)}{L(\chi_{1},1/2-iz)}-\frac{L^{\prime}(\chi_{2},1/2-iz)}{L(\chi_{2},1/2-iz)}$
$=\frac{1}{2}\log(N_{1}/N_{2})+\sum_{n\geq 2}\frac{\Lambda(n)(\chi_{1}(n)-\chi_{2}(n))}{\sqrt{n}}e^{2\pi i\tfrac{\log(n)}{2\pi}z}$
when ${\rm Im}\,z>1/2$, where $\Lambda(n)$ is the von Mongoldt function. It follows that $f=i\frac{A_{1}}{B_{1}}-i\frac{A_{2}}{B_{2}}$ is almost periodic in $\mathbb{C}^{+}+i/2$. The previous results show that the pair
$\mu$ $=2\pi\bigg(\sum_{L(\chi_{1},1/2+i\gamma_{1})=0}{\boldsymbol{\delta}}_{\gamma_{1}}-\sum_{L(\chi_{2},1/2+i\gamma_{2})=0}{\boldsymbol{\delta}}_{\gamma_{2}}\bigg)$
$a$ $=\log(N_{1}/N_{2}){\boldsymbol{1}}_{0}+\sum_{n\geq 2}\frac{\Lambda(n)}{\sqrt{n}}\bigg[(\chi_{1}(n)-\chi_{2}(n)){\boldsymbol{1}}_{\tfrac{\log(n)}{2\pi}}+(\overline{\chi_{1}(n)}-\overline{\chi_{2}(n)}){\boldsymbol{1}}_{-\tfrac{\log(n)}{2\pi}}\bigg]$
is a $\mathrm{FS}$-pair (above, ${\boldsymbol{1}}_{c}(\lambda)=1$ if $\lambda=c$ and ${\boldsymbol{1}}_{c}(\lambda)=0$ otherwise).
In what follows we say that an entire function $E$ is of Pólya class if $E$ has no zeros in $\mathbb{C}^{+}$, $|E^{*}/E|\leq 1$ in $\mathbb{C}^{+}$ and $y\mapsto|E(x+iy)|$ is nondecreasing for $y>0$. A classical result [13, Thm. 7] states that a function $E$ is of Pólya class if and only if
$E(z)=E^{(r)}(0)(z^{r}/r!)e^{-dz^{2}-ibz}\prod_{n}(1-z/\overline{z_{n}})e^{z{\rm Re}\,1/z_{n}},$ (9)
@@
The Pólya class can be also defined as functions that are uniform limits in compact sets of polynomials with no zeros in $\mathbb{C}^{+}$ (see [13, Problem 12]). Moreover, any function of Pólya class has order at most $2$ ([13, Problem 10]).
@@
Let $(\mu,a)$ be an $\mathrm{FS}$-pair. Assume that:
@@
    There is $b>0$ such that $\operatorname{\mathrm{supp}}(a)\cap(0,b)=\emptyset$ and $a(\cdot)$ has exponential growth.
Then there exists an entire function E=AiBE=A-iB of Hermite-Biehler class and of finite exponential type such that $E,E^{*}\in\mathrm{AP}(\mathbb{C}^{+})$ (and so $A/B\in\mathrm{AP}(\mathbb{C}^{+})$), ${\rm spec}(E)$ is bounded and (7) holds. If in addition $\operatorname{\mathrm{supp}}(a)$ is locally finite, then $E$ is a trigonometric polynomial and Remark 4 applies.
@@
Step 1. We first apply Theorem 3 to conclude there is an Hermite-Biehler function E=AiBE=A-iB such that $A/B\in\mathrm{AP}(\mathbb{C}^{+})$ and (7) holds. We can assume that $E$ has no real zeros as these can be removed (with a Weierstrass product) without altering $A/B$ or the fact that $E$ is Hermite-Biehler. In particular, $\operatorname{\mathrm{supp}}(\mu)=\{\gamma\in\mathbb{R}:\varphi(\gamma)\equiv 0\,({\rm mod}\,\pi)\}=\text{Zeros}(B)$, where $\varphi$ is the phase function associated with $E$, and these zeros are all simple. The zeros of $A$ are also simple and interlace those of $B$. Since $\mu$ is bounded from below on its support and $a(\cdot)$ has exponential growth, we can apply Remark 2 to obtain
@@
for some $0<c<1$, whenever $B(\gamma)=0$. In particular, we have $\sum_{B(\gamma)=0}\frac{1}{1+\gamma^{2}}<\infty$ and, since the zeros of $A$ interlace those of $B$, we also have $\sum_{A(s)=0}\frac{1}{1+s^{2}}<\infty$. We claim that we can assume $E$ is has order at most $1$. Indeed, define the canonical products
@@
for some $p,q\in\{0,1\}$, where $b\in\mathbb{R}$ is chosen so that $(A(z)/B(z))/(\mathcal{A}(z)/\mathcal{B}(z))=1+O(z)$ for $z\to 0$. Hence both $\mathcal{A}$ and $\mathcal{B}$ have order at most $1$ by a classical result of Borel [32, Thm. 6, p. 16]. Since ${\rm Re}\,iA/B\geq 0$ in $\mathbb{C}^{+}$ and, a straitfoward computation, shows that ${\rm Re}\,i\mathcal{A}/\mathcal{B}\geq 0$ in $\mathbb{C}^{+}$, we conclude that both $A/B$ and $\mathcal{A}/\mathcal{B}$ are of bounded type and so $F=(A/B)/(\mathcal{A}/\mathcal{B})$ also is and has no zeros. Since $F=F^{*}$, a classical result of Krein [25] (see also [33, Thm. 1, p. 115]) shows that $F$ is of exponential type, and so $F(z)=e^{hz}$, for some $h\in\mathbb{R}$. Since $F$ is of bounded type we have $h=0$. We can then define $\widetilde{E}=\mathcal{A}-i\mathcal{B}$ to finish the claim.
Step 2. Let $\{\overline{z}_{n}=x_{n}-iy_{n}\}$ be the zeros of $E$ with $y_{n}>0$. We claim $\sup_{n}y_{n}<\infty$ and that $E$ is of Póyla class. Indeed, since $f=iA/B\in\mathrm{AP}(\mathbb{C}^{+})$, by Remark 3 we have $\Theta=E^{*}/E=(1-f)/(1+f)\in\mathrm{AP}(\mathbb{C}^{+})$. Since ${\rm spec}(f)\subset[0,\infty)$, a simple computation using Lemma 12(i) shows $\mathbb{E}\Theta(0)=\frac{1-\mathbb{E}f(0)}{1+\mathbb{E}f(0)}$. However, $2{\rm Re}\,\mathbb{E}f(0)=a(0)>0$ (because $\mu\geq 0$), thus $|\mathbb{E}\Theta(0)|<1$. Assume now $\mathbb{E}\Theta(0)\neq 0$. Then Lemma 12(i) shows straightforwardly that $\Theta(x+iy)$ has no zeros for large $y>0$. Assume otherwise that $\mathbb{E}\Theta(0)=0$, that is, $\mathbb{E}f(0)=1$. Since $a(\lambda)=\mathbb{E}f(\lambda)$ for $\lambda>0$ we deduce that ${\rm spec}(f)\subset\{0\}\cup[b,\infty)$ where $b=\inf\{\lambda>0:a(\lambda)\neq 0\}>0$ by assumption. Since $f$ has no zeros in $\mathbb{C}^{+}$, we can then apply Lemma 12(ii) to conclude that $b\in{\rm spec}(f)$. Writing f=12gf=1-2g, with ${\rm spec}(g)\subset[b,\infty)$ and $\mathbb{E}g(b)\neq 0$, we can apply Lemma 12(i) to conclude that $|g(x+iy)|<1$ for large $y>0$, and so
$\Theta(z)=\frac{g(z)}{1-g(z)}=\sum_{n\geq 1}g(z)^{n},$
@@
Enumerate the $\text{Zeros}(B)=\{\gamma_{n}\}$ and let $c>0$ be so small that $\gamma_{n+1}-\gamma_{n}\geq c$ for all $n$. Let $0<\varepsilon<c/10$ to be chosen later. Observe first that if ${\rm dist}(x,\{\gamma_{n}\})=|x-\gamma|\geq\varepsilon$ (with $\gamma=\gamma_{n_{0}}$) then $|x-\gamma_{n_{0}+k}|\geq\varepsilon+|k|c/2$, and so